[ present & archived diaries ][ Home ]
- 19 January
Start up dynasonde again for EISCAT CP-1 experiments, running every 6 minutes.
- 27 January
Some data gaps due to hard disk full.
- 5 March
I am back in Tromsø after 3 months in Lindau
- 7 March
Found HV fuse blown again (last time was 30 October 1997). Replaced it
and all seems fine. Continue with B-mode, every 12 minutes.
- 23 March
Start IDI1.sch which is a B-mode (8p) followed by a P-mode repeating every
6 minutes. The P-mode is a single frequency (3.15 MHz) chosen to have low
interference but away from the PRE/MF radar at 2.8 MHz for IDI (Imaging
Doppler Interferometry). This will be analysed at BAS. Raw P-mode data
being recorded at the rate of 12 CD-ROM a day. This continues at least
until midnight on 27 March since CP2 and CP-6 are running on the UHF and
VHF radars simultaneously.
- 25 March
Existing HPTX attenuation value in BTIK15.CFG was not set so the default
of 12dB has been used. I changed the value in BTIK15.CFG to 10dB at 15:51
UT. The Pmode used, and continues to use 6dB.
- 26 March
Changed BTIK15.CFG HPTX value to 9dB (from 10) and IDI-T2.CFG to 5dB from
6 (at an unspecified time but afternoon). For the Pmode at 3.15MHz the
forward power reading is now 6 kW and reverse power is <0.1kW.
- 27 March
First ski tour of inspection of receiving antennas. They are on average
about 1.5m above the snow. No damage to antennas except a broken wire connecting
the north half of the North dipole to the pre-amp. Made a "hopefully
touching" connection at 13:45 UT. Back to main building to get tools
so that the "proper" repair was made between 15:31 and 15:33
UT. This will affect both B-modes and P-mode soundings this week. Who knows
how long the wire has been broken ?
- 21 Sept '98
about 1130 UT a power glitch (??) caused HV and Filament to switch off,
and HV fuse blown. After new fuse, HV can switch on but Iavg>500mA (off scale)
at 6.2 kV. HV cuts out but fuse still OK. Replaced keying transistor,
also put in the new fan at rear of HPTx (The fan had to be put in
backwards compared to the old one). However HV keeps cutting out !
Much extended trouble-shooting (see hand-written log book for
details). Thinking that D2-7 are the cause they were all replaced
with new ones. The other HPTx fan also replaced.
- 23 Sept '98
Finally, removing C8 (15uF, 1kV), the HV stays on !
- 24 Sept '98
Replace C8 with a 2uF, 2500VDC oil capacitor from EISCAT.
- 1-2 Nov '98
Upgraded the Dynasonde PC hardware from a 486, 33MHz 16MB and 430MB
machine to a 133MHz Pentium with 130 MB RAM and 3.2 GB hard disk.
Simultaneously went from DOS 6.0 to 6.22, which meant putting /Y
switch in the COPY and MOVE commands to force file overwriting.
Upgrade successful after disabling LPT1 (378 and IRQ7) which caused
a conflict with the PC clock, both with and without CRAMCLOCK
- 6 Nov
Upgraded the network software on the dynasonde PC to use MS-DOS
Network client V3.0 instead of the windows for workgroups 3.0 stuff.
This has the effect that an attempted access to a non-existant
network drive times out after a few seconds. Earlier was infinitely
long causing the sounding schedule to stop.
I also put a test for the network drive in the EXECUTE statement.
- 17-18 Nov
Ran IDI mode plus B-modes for the the Leonids meteor shower.
See results processed by Owen Jones (BAS). Some good velocities
near the end.
[ Home ] This page is maintained by Mike,
who is solely responsible for the content, or lack thereof :-/ Last
updated Wednesday, 3-Mar-99 15:24:53 MET DST